### v1

# Structural Change, Employment, and Inequality in Europe: an Economic Complexity Approach

Structural Change, Employment, and Inequality in Europe: an Economic Complexity Approach
Bernardo Caldarola Corresponding author. Email: caldarola@merit.unu.edu Enrico Fermi Research Center, Rome, Italy UNU-MERIT, Maastricht, the Netherlands School of Business and Economics, Maastricht University, Maastricht, the Netherlands Dario Mazzilli Enrico Fermi Research Center, Rome, Italy Aurelio Patelli Enrico Fermi Research Center, Rome, Italy Angelica Sbardella Enrico Fermi Research Center, Rome, Italy
Structural change consists of industrial diversification towards more productive, knowledge-intensive activities. However, changes in the productive structure bear inherent links with job creation and income distribution. In this paper, we investigate the consequences of structural change – defined in terms of labour shifts towards more complex industries – on employment growth, wage inequality and functional distribution of income. The analysis is conducted for European countries using data on disaggregated industrial employment shares over the period 2010 – 2018. First, we identify patterns of industrial specialisation by validating a country-industry industrial employment matrix using a bipartite weighted configuration model (BiWCM). Secondly, we introduce a country-level measure of labour-weighted Fitness, which can be decomposed in such a way as to isolate a component that identifies the movement of labour towards more complex industries – the structural change component. Thirdly, we link structural change to i) employment growth, ii) wage inequality, and iii) labour share of the economy. The results indicate that our structural change measure is associated negatively with employment growth. However, it is also associated with lower income inequality. As countries move to more complex industries, they drop the least complex ones, so the (low-paid) jobs in the least complex sectors disappear. Finally, structural change predicts a higher labour ratio of the economy; however, this is likely to be due to the increase in salaries rather than by job creation.
Keywords: Structural Change, Economic Complexity, Inequality, Employment JEL Codes: D63, E24, J31, O11, O52
1 Introduction
Structural change can be defined by the reallocation of economic activity across bread sectors of the economy (Lewis 1954, Syrquin 1988). The shift of the relative importance of industrial employment or output towards increasingly more productive activities that defines the process of structural change is also an engine of long-term, sustained economic growth (Herrendorf et al. 2014, 2022). However, changes in the productive structure are also inherently associated with the distribution of income, as highlighted in the seminal work by Kuznets (1955). For instance, when countries upgrade their productive structure, they also move to sectors with higher knowledge intensity, and lower labour requirements. This may have consequences on the wage and functional distribution of income, as well on the creation of new jobs and the destruction of old ones.
In this paper, we use the case of Europe to study the relationship between the movement of labour to increasingly knowledge intensive activities – defined here as structural change – with employment growth, wage inequalities, and the functional distribution of income between labour and capital.
As a matter of fact, in the last decades within country inequality has been growing in Europe and in high-income countries. At the same time, these countries have de-industrialised during the last wave of ICT-led globalisation, leading to the substitution of manufacturing with jobs in complex, knowledge intensive industries – such as business services – contributing to increasing inequalities (Antonelli & Tubiana 2020). The mechanism underlying this trend has been explored further in the empirical literature. The Skill-Biased Technological Change (SBTC) literature (Autor et al. 2003, Autor 2022, Acemoglu & Restrepo 2022) has examined the impact of the main driver of structural change – technological change – on employment and distribution of earnings, suggesting that waves of technological change tend to complement only more educated workers in knowledge intensive industries, leaving less educated workers behind.
Moreover, structural change bears a strong linkage with employment creation. As economies transform in their productive structures, they create winners and losers (Schumpeter 1934). The current wave of technological change may not reduce total employment (Dauth et al. 2021), but induces labour market adjustments at the disadvantage of low-skilled workers, whose share over total workers is reduced (Graetz & Michaels 2018). In the case of other technologies, such as ICT, these may have been the culprits of jobless growth (Brynjolfsson & McAfee 2011, Frey & Osborne 2017) and contributed to the shrinking of the number of middle-skill jobs (Goos et al. 2009, 2014).
In order to identify the degree of knowledge intensity of European industries we employ an Economic Complexity (EC) perspective. Complexity metrics aim at quantifying the capability structure of countries by looking at their productive basket, preserving information on their specialisation and diversification patterns rather than by simply aggregating their economic output (Hausmann & Hidalgo 2009, Tacchella et al. 2012). By the same token, EC measures allow to identify the capability (knowledge) intensity (Lo Turco & Maggioni 2022) of individual products, industries, or technologies by considering their ubiquity – a proxy for the diffusion of the capabilities that such activities require – weighted by the capability endowment of the countries specialised in such activities – i.e., the countries’ Fitness. Given the strong link of EC metrics with economic diversification, and in particular of the Economic Fitness and Complexity (EFC) approach (Tacchella et al. 2012), they have proved to be a useful tool to describe a long-term dynamic such a structural change (Freire 2021, Castañeda et al. 2022, McNerney et al. 2023), being also highly predictive of GDP growth (Tacchella et al. 2018). Following the analogy between structural change and complexity, the latter can be used to explain the inequality trends that result of changes in the sectoral composition of the economy. The literature on complexity and inequalities (Sbardella et al. 2017, Hartmann et al. 2017, Hartmann & Pinheiro 2022) indicates that higher complexity of the productive structure is associated to lower inequalities at the country level. It is also well known that knowledge concentrates in space (Balland et al. 2020, 2022), turning geography into a useful lens through income distribution can be analysed. In fact, at the regional level the sign of this relationship flips (Sbardella et al. 2017), as only regions with the strongest knowledge basis are able to diversify into highly complex industries, leaving laggards behind (Pinheiro et al. 2022). Looking at the relationship between complexity and employment, Adam et al. (2023) finds that higher complexity is linked to job creation. Where existing jobs are made redundant, it has been shown that re-employment is easier in presence of related industrial variety at the regional level (Hane-Weijman et al. 2018).
One important feature of EC methods is that these allow to quantify the capability requirements of finely disaggregated industries, offering a granular description of the process of structural change both across and within broad sectors of the economy. In order to offer an encompassing assessment of the complexity of European industries and of the labour flows across them, we depart by traditional approach of the EC literature – which infers the productive structure of countries using export data – by using data on industrial employment in European countries between 2010 and 2019. First, we consider the use of employment data to measure industrial Complexity and countries’ employment Fitness as this allows to include also non-tradeable activities, most of which are in the service sector. This is particularly relevant considering that European economies have heavily shifted their productive structure towards the tertiary industries along with the process of globalisation. Secondly, the focus on employment represents a crucial element in the conceptualisation of structural change as the change in relative shares of the economy. Using employment data to measure Economic Fitness and Complexity respectively at the country and industry level resonates with this conceptual framework, and allows to outline the structure of capabilities of labour as a productive input.
In the present study, the analysis of the relationship between structural change, inequality, and employment creation is conducted for European countries over the period 2010 -– 2018. The empirical approach relies on information on European industrial employment provided by Eurostat’s Structural Business Survey data, measured at a fine level of disaggregation (NACE 4 digits) across manufacturing and service industries. This information is matched with data on the distribution of wages at the country level, provided by the ILO, and with data on the labour share of the economy and employment growth, obtained by ARDECO. Using this newly constructed dataset, we start by identifying patterns of industrial specialisation by validating the country-industry industrial employment matrix using a bipartite weighted configuration model (BiWCM), overcoming some limitations imposed by the Balassa (1965) method. Secondly, we introduce a measure of labour-weighted Fitness, which sums up complexity of industries weighted by their employment share. This measure can be decomposed in such a way as to identify the contribution to changes in labour-weighted Fitness coming from the movement of labour towards more complex industries. We identify such component as the one linked to structural change from a labour viewpoint – the structural change component. Thirdly, we link the structural change component to a number of economic outcomes at the country level: i) employment growth, ii) wage inequality, and iii) functional distribution of income (labour share of the economy). In this way, we analyse the distributional consequences of structural transformation – framed as the movement of labour to more complex industries – and its job-creating potential.
An analysis of the specialisation patterns in European countries, conducted by inspecting the significant links of the country-industry specialisation matrix generated by the BiWCM, reveals that the most diversified countries have progressively moved away from specialisation in low-complexity, labour-intensive industries. Moreover, our empirical analysis explores the association between the structural change component of our labour-weighted Fitness measure -– derived from the movement of labour towards more complex industries -– to employment growth, wage inequality, and the functional distribution of income (labour share of the economy). The preliminary results of OLS panel regressions with country and time fixed effects indicate that our structural change measure is associated negatively with employment growth, corroborating the evidence that highly complex industries have lower labour requirements. However, it is also associated with lower income inequality – measured in terms of the ratio of average wages in the ninth and first deciles of the wage distribution. As countries move to more complex industries, they drop the least complex ones, so the (low-paid) jobs in the least complex sectors disappear, making the first decile of the salary distribution go up. Finally, structural change predicts a higher labour ratio of the economy; however, this is likely to be due to the increase in salaries in highly complex industries rather than by job creation. Very interestingly, the structural change component is the only one to be significant in the regressions, while the increase in within-sector complexity, and the labour-weighted Fitness of countries do not explain our outcome variables. Results are robust to a battery of robustness checks, including leave-one-out regressions and a specification where the structural change component of the labour-weighted Fitness is replaced by a similar component obtained decomposing a labour-weighted entropy measure.
The remainder of this paper is organised as follows. In Section 2 we delve into the Economic Fitness and Complexity framework, and we introduce a new measure – the Labour-Weighted Fitness (LWF) the variation of which can be decomposed to extract a structural change component. In Section 3 we outlined the data sources used in the analysis, which in turn is described in the following Section (4). Section 5 illustrates the preliminary results of the analysis, and Section 6 concludes.
2 Analytical Framework
In this section, we will outline the analytical framework used to derive a measure of structural change based on the reallocation of labour to complex industries. First, we will define countries’ Fitness and industrial complexity according to the EFC framework (Tacchella et al. 2012). Second, we introduce a new measure – the Labour Weighted Fitness – from which we derive a structural change component.
2.1 Structural change, Fitness and Complexity
Economic Complexity is a framework that builds on earlier evolutionary and institutional literature (Cimoli & Dosi 1995, Hirschman 1958, Teece et al. 1994) to tackle the complexity of economic systems. It describes the economy as an evolutionary process of globally interconnected ecosystems. The main recent advance with respect to the earlier literature is the use of newly developed network science and other methods to investigate complex and dynamical systems (Hausmann & Klinger 2006, Hausmann & Hidalgo 2009, Tacchella et al. 2012) to separate the random noise from the underlying signal. The Economic Complexity framework shifts the focus from aggregate quantities (What is the GDP of the country?) to a more disaggregated view (In which industrial sectors do countries specialise?), with the aim of providing information that is complementary to more traditional analysis. This shift creates the opportunity to create a consistent framework that allows several cross-cutting themes to be addressed and provides a quantitative answer to several policy relevant questions that could otherwise only be answered qualitatively or by ad hoc metrics.
In particular, EC methods have contributed “to the study of structural change by addressing questions related to how the diversification of countries and the patterns of trade specialization affect and are affected by economic growth, and how diversification and trade specialization change over time, and the determinants of such change” (Freire 2021, p 214), providing useful tools to study disaggregated aspects of structural change.
In order to measure the Fitness of countries, and more importantly for our purposes, the Complexity of European industries, we rely on the Fitness and Complexity algorithm proposed by Tacchella et al. (2012):
$\widetilde{F}_{c}^{(n)}=\sum_{i}M_{ci}Q_{i}^{(n-1)}&&F_{c}^{(n)}=\frac{\widetilde{F}_{c}^{(n)}}{\langle\widetilde{F}_{c}^{(n)}\rangle_{g}}\\ \\ \widetilde{Q}_{i}^{(n)}=\frac{1}{\sum_{c}M_{ci}\frac{1}{F_{c}^{(n)}}}&&Q_{i}^{(n)}=\frac{\widetilde{Q}_{i}^{(n)}}{\langle\widetilde{Q}_{i}^{(n)}\rangle_{a}}\end{array}\right.$ (1)
with $F^{(0)}_{c}=1$ and $Q^{(0)}_{i}=1$.
The algorithm is fed with a country-industry employment matrix constructed by cross-tabulating industrial employment data across European countries, adequately binarised [^1] The binary specialisation matrix – which identifies industries in which countries are specialised – is obtained using an entropy maximising null model (the Bipartite Weighted Configuration Model) and implementing a method to ensure the comparability of fitness and complexity values over time. A more in-depth explanation of the construction of the matrix are explained in section 4.1, from which we obtain the values Fitness (first row of equation 1) and Complexity (second row of equation 1). In practical terms, the Fitness $F$ of country $c$ at each time period corresponds to the sum of the Complexities $Q$ of the industries $i$ in which country $c$ holds a comparative advantage, the latter being defined by the row sums of the binary specialisation matrix $M_{c,i}$. In turn, the complexity $Q$ of industry $i$ corresponds to the inverse of its ubiquity (the column sums of $M_{c,i}$) weighted by the inverse of the Fitness of the countries which have that industry in their productive portfolio. In short, more complex industries are the least ubiquitous ones – proxying for the rareness of the capabilities required by those – considering also how "fit" are the countries that hold a specialisation in that industry. The algorithm is iterated $n$ times, until it converges. At each step, Fitness and Complexity values are normalised by dividing them by the respective average values.
A feature that makes Fitness our preferred measure of knowledge intensity of the productive structure is that it considers the cumulative nature of capabilities, giving importance to the diversification of countries (weighted by the complexity of their industries). This feature is absent in competing methods, such as the Economic Complexity Index (ECI) (Hausmann & Hidalgo 2009), which has been shown to measure country fitness as the average complexity of the economic activities in which a country holds a revealed comparative advantage (Cristelli et al. 2013).
2.2 Labour-Weighted Fitness and decomposition
The starting point to construct a measure of structural change that captures the reallocation of labour across industries is to manipulate the Fitness equation (the first row of equation 1) in such a way as to weight the sum of the complexities $Q_{i}$ by their labour share. Thus, the Labour-Weighted Fitness can be expressed as:
$LWF_{t}=\sum_{i=1}^{n}\theta_{i,t}Q_{i,t}$ (2)
Where $\theta_{i,t}$ is the labour share of industry $i$ at time $t$ for any given country. The resulting LWF is therefore an index of the Fitness of countries, where more weight is given to industries that are not only complex, but which also absorb a larger employment share, indicating the presence of knowledge-intensive capabilities across a larger portion of the workforce.
To extract our measure of structural change, we now propose a structural decomposition of LWF inspired by the productivity decompositions adopted in the structuralist literature (McMillan et al. 2014, 2017, de Vries et al. 2015), which originates from the original work of Fabricant (1942). We decompose $\Delta LWF_{t}$ – the variation of LWF between time $t$ and $t-k$, with $k\leq T$ – in within and between sector components, as follows:
$\Delta LWF_{t}=LWF_{t}-LWF_{t-k}=$ (3)
$(=)\ \sum_{i=1}^{n}\theta_{i,t}Q_{i,t}-\sum_{i=1}^{n}\theta_{i,t-k}Q_{i,t-k}=$ (4)
$(=)\ \sum_{i=1}^{n}\theta_{i,t}Q_{i,t}-\sum_{i=1}^{n}\theta_{i,t-k}Q_{i,t-k}\color[rgb]{0,0.6,0}{+\sum_{i=1}^{n}\theta_{i,t-k}Q_{i,t}-\sum_{i=1}^{n}\theta_{i,t-k}Q_{i,t}}=$ (5)
$(=)\ \underbrace{\sum_{i=1}^{n}\theta_{i,t-k}\Delta Q_{it}}_{\text{WITHIN TERM}}+\underbrace{\sum_{i=1}^{n}Q_{it}\Delta\theta_{i,t}}_{\text{BETWEEN TERM}}$ (6)
After manipulation, Equation 3 yields two distinct terms, described in Equation 6:
  • $\sum_{i=1}^{n}\theta_{i,t-k}\Delta Q_{it}$ is the within-sector term of structural change (the contribution of increasing complexity at the sectoral level to the change in labour-weighted Fitness)
  • $\sum_{i=1}^{n}Q_{it}\Delta\theta_{i,t}$ is the between-sector term of structural change (the contribution of labour shifts from low to high complexity sectors to the change in labour-weighted Fitness)
The between component, given by $\sum_{i=1}^{n}Q_{it}\Delta\theta_{i,t}$, identifies the measure of structural change of interest. This will be used to explain the observed trends in wage inequality, employment growth, and functional distribution of income.
3 Data
3.1 EUROSTAT’s Structural Business Survey
The main data source used in this paper is the Eurostat Structural Business Statistics (SBS). The SBS describe the structure and performance of industrial sectors in the in European countries. They cover the ‘business economy’ (NACE Rev. 2 sections B to N and division 95) which includes: i) Industry; ii) Construction; iii) Distributive trades; and iv) Services. The data on industries is aggregated at the 4 digit level of the NACE classification, for a total of 218 NACE industries.
SBS are based upon data for enterprises, which is obtained for national statistical offices and aggregated at the industry level based on the enterprise’s principal activity. Businesses in financial services (NACE Rev. 2 Section K) are excluded because of their specific characteristics and the unavailability of information in this industry, often due to confidentiality issues. In addition, SBS does not cover agriculture, forestry and fishing (NACE A), public administration (NACE O) and non-market services such as education (NACE P) and health (NACE Q).
As the analysis conducted here relies on patterns of industrial employment, we use the SBS data to describe the employment share of 4 digit NACE industries. To do so, we extract the SBS indicator that quantifies the number of persons employed [^2] Following the Eurostat’s SBS definition, this indicator ”is defined as the total number of persons who work in the observation unit (inclusive of working proprietors, partners working regularly in the unit and unpaid family workers), as well as persons who work outside the unit who belong to it and are paid by it (e.g. sales representatives, delivery personnel, repair and maintenance teams). It excludes manpower supplied to the unit by other enterprises, persons carrying out repair and maintenance work in the enquiry unit on behalf of other enterprises, as well as those on compulsory military service.” Source: https://ec.europa.eu/eurostat/cache/metadata/en/sbs_esms.htm#meta_update1663843213928 in each industry. In order to maximise data availability, we restrict the analysis to the 27 countries that were part of the European Union until 2018, observed between 2010 and 2018.
One limitation of the SBS data is that it presents a large number of missing data across countries, industries and years. Given that the the construction of the indicators used in the analysis (such as the labour-weighted fitness illustrated in section 2.2) requires full series of industrial employment shares across countries and years, we have resorted to a data reconstruction strategy, based on linear interpolation and extrapolation at constant values. This strategy has proved to be the best performing in terms of its ability to correctly predict missing values. A detailed description of the data reconstruction strategy, and a comparison between competing methods, is presented in appendix A.
3.2 ARDECO
The Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy (ARDECO) [^3] The ARDECO dataset is available, free of charge, from the European Commission’s website: https://knowledge4policy.ec.europa.eu/territorial/ardeco-database_en, is a dataset curated and maintained by the Commission’s Joint Research Centre. It includes information on European countries and regions at different levels of disaggregation. While the dataset is based on the Eurostat’s SBS illustrated in the previous section, the data is complemented – where needed – with national and international sources. The data covers different topics, such as population, employment, labour costs, domestic product and capital formation. For the purpose of the analysis carried out in this paper, indicators on population, GDP per capita (constant prices), employment growth, compensation of employees (at constant prices) and gross value added (at constant prices) have been selected. The last two indicators are used to construct a measure of labour share of the economy, computed as the ratio of employee’s aggregate compensation over gross value added, in each country and year.
3.3 ILO Statistics on Labour Income and Inequality
The data on wage inequality is obtained from the International Labour Organisation’s Statistics on Labour Income and Inequality. We focus on data regarding the labour income distribution [^4] Available at: https://www.ilo.org/shinyapps/bulkexplorer46/?id=LAP_2LID_QTL_RT_A, which despite being based on modelled estimates, it provides information on average wages at all deciles of the labour income distribution across the countries in the sample, for the entire time period covered (2010 – 2018). In particular, we use information on labour income at the ninth, fifth and first deciles to produce wage ratios between ninth and first, fifth and first, and ninth and fifth deciles to capture the degree of inequality in labour income at different points of the distribution.
3.4 World Development Indicators
The data described above is complemented with additional information extracted from the World Bank’s World Development Indicators to further characterise the profile of the countries in our sample, and to mitigate the omitted variable bias. As we are focusing on time-varying characteristics of countries that have somehow to do with their productive structure and technological capabilities, we use information on the share of exports and share of R&D investments over total GDP.
4 Methods
4.1 Specialisation matrix using BiWCM and dummy fitness
Economic Complexity applications frequently involve the need to determine the areas in which individuals or entities possess a comparative advantage. This categorization is established using a tagging scheme, which is computed using the Revealed Comparative Advantage (RCA) index, originally formulated by Balassa (Balassa 1965). In the context of our current study, the entities being assessed are countries, and the areas of specialization are represented by the number of workers in various industries. The RCA index quantifies this specialization by calculating the ratio of a country’s share of workers in a particular industry to the global share of that same sector. To express this mathematically, the RCA is:
$[RCA]_{c,i}=\frac{L_{c,i}}{\sum_{i^{\prime}}L_{c,i^{\prime}}}\left/\frac{\sum_{c^{\prime}}L_{c^{\prime},i}}{\sum_{c^{\prime},i^{\prime}}L_{c^{\prime},i^{\prime}}}\right.=\frac{L_{c,i}}{E[L]_{c,i}},\quad E[L]_{c,i}=\frac{(\sum_{i^{\prime}}L_{c,i^{\prime}})(\sum_{c^{\prime}}L_{c^{\prime},i})}{\sum_{c^{\prime},i^{\prime}}L_{c^{\prime},i^{\prime}}}$ (7)
$M_{c,i}$ is the matrix used for example in eq. (1) and is the binary filter of the RCA associating 1 when the RCA value is above the unity, otherwise it is zero.
The primary rationale for utilizing the RCA index is to establish a measure that isn’t contingent on the relative strength of diverse countries and industries, which are inherently incomparable in terms of size. Balassa’s RCA, specifically, contrasts the actual labour value of a country with an anticipated value derived under the assumption of equalising country and sector sizes (Vollrath 1991, Krantz et al. 2018). In this interpretation, RCA is a ratio between the actual number of workers and the expected output of an average country of the same size, given the knowledge of only the total labour strength, as made explicit in the second part of eq. (7). However, it’s essential to acknowledge that RCA carries inherent bias due to the heterogeneity of the country and industrial strength, resulting in correlations in the estimation of average country values, as discussed in (Bruno et al. 2023). In statistical terms, RCA doesn’t satisfy the maximum likelihood condition, whereas entropy-based null models provide unbiased estimations (Garlaschelli & Loffredo 2008).
In this study, we employ the Inferred Comparative Advantage (ICA) to construct the $M_{c,i}$ matrix, comparing observed labour force with estimates from an entropy-based unbiased null-model which maintains fixed country-wide total labor and total sector strength. The null-model is known as the Bipartite Weighted Configuration Model (BiWCM) (Bruno et al. 2023). We adopt a similar methodology to RCA, adjusting for the accurate definition of the average country $E[L]_{c,i}$ by computing it using the BiWCM, such that the interpretation of the results is straightforward in the Economic Complexity framework. However, incorporating a null model through the BiWCM allows for more precise identification of comparative advantages with statistical significance. For instance, figure 1 illustrates the p-values of empirical values $L_{c,i}$ using a colour scale ranging from blue (indicating zero) to red (indicating one). Before feeding the ICA ($M_{c,i}$) matrix described in figure 1 into the EFC algorithm outlined in equation (1), we add an additional step to ensure comparability of Fitness and Complexity values across years.
The EFC algorithm computes a fixed point to determine the actual values of the Fitness and Complexity indices for each year independently. Consequently, in principle, it is not possible to compare the index values across different years due to the dynamic changes in the bipartite networks. To address this issue, we employ a technique, the "dummy trick", as elaborated in (Mazzilli et al. 2024). This involves augmenting the $M_{c,i}$ matrix with an additional row, referred to as the "dummy" row, which represents a country specialised in all industries. Thanks to the scale invariance property of the algorithm, we assign a constant Fitness value of 1 to this dummy country for every year. This approach establishes a fixed reference scale over time, enabling the comparison of Fitness and Complexity values across different years.
4.2 Econometric framework
It is now possible to compute the Industrial Complexity Index $Q_{i}$ used to construct our measure of Labour-Weighted Fitness, as described in equation 2, which in turn is decomposed into a within and a between term (equation 6). In particular, as mentioned in Section 2.2, we are interested in the decomposition term that captures the shift of labour shares across industries, that is, the between term. We use this term to explore its correlation with employment growth, wage inequality and the labour share of the economy, using the following specification:
$Y_{c,t}=\alpha+\beta_{1}between_{c,t}+\beta_{2}within_{c,t}+\beta_{3}\Delta LWF_{c,t}+\beta_{4}X_{c,t}+\sigma_{c}+\tau_{t}+\epsilon_{c,t}$ (8)
Where $between_{c,t}$, $within_{c,t}$ and $\Delta LWF_{c,t}$ are respectively the between term, within term, and sum of the previous two, as described by equation 6; in all cases, these quantities are computed for $k=1$. $Y_{c,t}$ is a placeholder for measures of employment growth, wage inequality and labour share of the economy. Employment growth is computed as the share of employed, working age population in each country at time $t$ over the same share of employment in the previous time period, as obtained from the ARDECO dataset:
$g_{c,t}=\frac{emp_{c,t}}{emp_{c,t-1}}$ (9)
Wage inequality is constructed using the ILO Statistics on Labour Income and Inequality described in Section 3.3. In order to capture the effect of structural change on the distribution of income, we take three different measures of wage inequality: 9th to 1st percentile, 9th to 5th percentile, and 5th to 1st percentile:
$ineq_{c,t}^{p^{th}/q^{th}}=\frac{wage_{c,t}^{pth}}{{wage_{c,t}^{qth}}}$ (10)
With the superscripts $p$ and $q$ indicating either the 9th, 5th or 1st percentile. Finally, we use the ARDECO data to compute a measure of the labour share of the economy in each country and year, taking the ratio of the total wages paid to workers over the total value added, all measured at constant prices (2015):
$labshare_{c,t}=\frac{wage_{c,t}}{VA_{c,t}}$ (11)
Finally, $X_{c,t}$ is a vector of time varying controls at the country level, including GDP per capita, population (from ARDECO), share of R&\&&D investments over GDP, share of exports over GDP (from the World Development Indicators).
All equations are estimated using OLS country ($\sigma_{c}$) and time ($\tau_{t}$) fixed effects, with errors clustered by country and year. Preliminary results of the descriptive econometric model are provided in the next section, along with an illustration of the results produced by the construction of the specialisation matrix and structural decompositions.
5 Results
5.1 Descriptive results
We begin by commenting on the patterns exhibited by the specialisation matrix, which has been obtained by filtering the country-industry employment matrix for the year 2010 as described in Section 4.1. Figure 1 contains information on the statistical significance of individual country-industry linkages, with values going from more significant (red) to less significant (blue). The rows of the matrix represent countries, while the columns identify industries. Countries have been ordered from higher Fitness (top) to lower Fitness (bottom), and industries are sorted from the least (left) to the most (right) complex.
Figure 1: BIWCM Specialisation Matrix
At a first visual inspection, the matrix displays a pattern ascribable to the nested pattern (Bustos et al. 2012) that typically characterises country-product RCA matrices, filtered using the Balassa index (Patelli et al. 2023). Precisely, the distribution of the statistically significant linkages between countries and industries that describes countries’ diversification patterns has a triangular shape, which can be used as a broad approximation of the nested nature of the system being observed. While the concept of nestedness has originated in evolutionary biology, it has found vast application in the economic complexity literature: productive systems that exhibit a nested pattern reveal the presence of correlations within the system that cannot be accounted for using linear models or average approximations. In economic terms, nestedness indicates that industrial activities that are pursued by low-Fitness countries are also pursued by high-Fitness countries, which are however also more diversified.
In the specific case described in figure 1, the economic interpretation of nestedness holds only in part; as a matter of fact, the top left corner – where high-Fitness countries and low-Complexity industries are located – appears to be less populated that one would expected in a fully nested matrix. While a similar pattern emerges also in other systems – such as the one of scientific production (Patelli et al. 2023) – in this case we find the observed result consistent with the process of structural change. When it comes to the employment dimension of the industrial production, high-Fitness countries tend to have higher diversification and to develop a comparative advantage in capital- or knowledge-intensive industries, at the expensive of more labour-intensive (and low-complexity) ones. As a result, industrialising countries with a comparative advantage in labour intensive activities (often due to lower wages) tend to specialise in the latter industries, and to have lower diversification. In substance, the sample of European countries seems to be divided in two subsamples of countries at different stages of structural transformation, with a minority of more mature countries and a majority of industrialising countries. The dynamic of Fitness rankings shows clearly which countries belong to each group, with the more mature economies being Germany, France, Italy, United Kingdom and Spain, followed by emerging countries such as Poland and Czechia.
Figure 2: Fitness rankings: 2010-2018
We now move towards an analysis of the variation in LWF over the time period 2010–2018. Figure 3 shows the difference in LWF between 2010 and 2018, computed as per equation 6, with its respective between and within component. We shall recall that the former measures the contribution of structural change – defined here as the movement of labour to more complex industries – to the variation in LWF over the time period under consideration; while the latter picks up the change in LWF due to increasing complexity of the industries in which countries develop an (inferred) comparative advantage. Overall, figure 3 indicates that the between and within component are often anticorrelated. This can be read in economic terms by considering the reasonable assumption that as industries become more complex, they also become less labour intensive. The within component captures if the industries in which a country is specialised become more complex – for instance, as a result of technical change within the industry. By doing so, labour is shed by such complex industries, flowing into industries with lower complexity – a pattern associated with a between component of the opposite sign to the within component.
To have a clearer idea of what happened in European countries over the 2010–2018, figure 4 describes the same quantities as figure 3, but for two different time periods: 2010–2014 (left panel) and 2014–2018 (right panel). The figure indicates that the first time period the variation in LWF fitness was negative for most countries, mostly driven by the decreasing within-industry complexity; it must be noted that this period coincides with the aftermaths of the 2008 financial crises, which has severely compromised economic activity in Europe across the board. In the following period, however, European countries show a much better performance, with most countries increasing their LWF, in many case due to within industry complexity and between industry reallocation of labour. One interesting case is the one of Poland, which is an outlier in both period: the country shows a huge within industry growth of complexity in 2010–2014, which is however followed by a negative between industry reallocation of labour in the following time period – hypothetically because of the lower labour intensity of its productive structure that has shifted towards increasingly knowledge-intensive activities.
Figure 3: Labour-weighted Fitness decomposition: 2010-2018
Figure 4: Labour-weighted Fitness decomposition: 2010-2014 and 2014-2018
5.2 Structural change and employment growth, wage inequality, and functional distribution of income
In this section we report the results of the estimation of equation 8 using the different outcome variables for employment growth, wage inequality and functional distribution of income. It should borne in mind that, at the present stage, the results presented below aim exclusively at describe the correlation between structural change and the outcome variables, without any implication for their causal relationships.
Table 1 summarises the relationship between structural change ($between$) and employment growth. Looking at columns (1) and (4), the between component of the LWF structural decomposition shows a negative correlation with employment growth at the country level. This indicates that when countries experience a labour flow from low- to high-complexity industries, they also experience slower rates of employment growth. This finding seems to corroborate the hypothesis that high-complexity industries contribute less to aggregate employment, due to their higher levels of knowledge intensity, and lower labour-intensity, as they require fewer (and likely more skilled) workers. It is worth mentioning that neither the within component of the variation in LWF (column 2), nor the sum of both components (column 3) do not show any significant correlation with employment growth, highlighting the importance of structural change in explaining outcomes related to the creation of employment. The coefficient on the between component remains significant also when the within component is also included in the regression.
We now move to the relationship between structural change and wage inequality. First, we look at the link between structural change and the wage gap between top (9th decile) and bottom (1st decile) wages. The wage gaps appears to contract as a result of labour reallocation towards knowledge-intensive industries, as indicated by columns (1) and (4) in Table 2. However, it is not clear whether this result is driven by lower wages in the top of the distribution, or higher wages at the bottom. In order to further explore this, Tables 3 and 4 show respectively the results for the wage gap between 5th/1st and 9th/5th deciles. Interestingly, the wage gap between median and bottom salaries narrows (columns 1 and 4 in Table 3) as a result of structural change, while the ratio between salaries at the 9th and 1st decile doesn’t show any association with structural change (Table 4, columns 1 and 4). We conclude that the reduction in the wage gap between top and bottom wages results from bottom wages growing at a faster rate than top wages. In order to rationalise this result, it is worth going back to the specialisation matrix described in figure 1. As countries move to more complex industries, they drop the least complex ones, as indicated by the "emptier" top-left corner of the matrix. Assuming a positive correlation between complexity and average wages, it can be inferred that the reduction in wage inequality is driven by the fact htat the low-paid jobs in the least complex sectors disappear, making the 1st decile of the salary distribution go up. Although this conjecture is yet to be tested at the present stage of the research, it represents one of the next steps in the analysis conducted in this paper.
Finally, Table 5 displays the results of regressing labour share (the ratio between aggregate wages and value added in the economy) on the between component of LWF. Columns (1) and (4) show a positive relationship between the two, indicating that as workers move to more complex industries, that contributes positively to the increase of the labour component of production, which appears to become more intensive. In order to understand this results, we should rely once more on the assumption that high-complexity industries pay, on average, higher wages. Bearing in mind the diversification patterns displayed by figure 1, coupled with the results on employment growth in Table 1. If employment growth slows down as a result of structural change towards knowledge-intensive industries, the increase in the labour share of the economy is due to be driven by higher average wages rather than higher participation of the population to the workforce. This hypothesis will be tested in the future steps that we will take to carry on with the analysis.
As a final note on the preliminary results presented in this section, it is important to stress that the structural change component is the only one to be significant across the regression specifications summarised by equation 8. As a matter of fact, neither the within component nor the change in labour-weighted fitness explain any of the outcomes observed, indicating that framing structural change as the movement of labour across industries of different complexity – that is, knowledge intensity – can contribute significantly to explain the current trends in employment growth, wage inequality and functional distribution of income in European countries.
Table 1: Employment growth regressions
[…]
FE OLS
Model: (1) (2) (3) (4)
Variables
Between -7,313.359∗∗ -7,527.217∗∗
(3,229.347) (3,248.861)
Within 1,266.110 1,412.623
(1,121.540) (1,123.097)
$\Delta C$ 279.081
Population (log) -19.942 -19.817 -19.927 -19.785
(12.878) (13.098) (13.057) (12.968)
GDPpc (log) -6.179 -5.824 -6.046 -5.859
(3.563) (3.572) (3.556) (3.570)
[…]
(0.064) (0.063) (0.063) (0.063)
[…]
(0.761) (0.736) (0.756) (0.718)
Fixed-effects
Country Yes Yes Yes Yes
Year Yes Yes Yes Yes
Fit statistics
Observations 224 224 224 224
R2 0.526 0.522 0.520 0.528
Within R2 0.233 0.225 0.223 0.236
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
Table 2: Wage inequality regressions: ratio $9^{th}/1^{st}$ dec.
Dependent Variable: Ratio $9^{th}/1^{st}$ dec.
FE OLS
Model: (1) (2) (3) (4)
Variables
Between -8,201.860∗∗∗ -8,142.939∗∗∗
(2,884.048) (2,893.616)
Within -547.693 -389.195
(671.490) (744.401)
$\Delta C$1,372.345
GDPpc (log) -3.781 -3.831 -4.031 -3.870
(2.770) (2.863) (2.861) (2.778)
Population (log) -5.596 -5.675 -5.763 -5.640
(7.401) (7.552) (7.514) (7.424)
[…]
(0.024) (0.025) (0.024) (0.024)
[…]
(0.995) (1.000) (1.000) (1.002)
Fixed-effects
Country Yes Yes Yes Yes
Year Yes Yes Yes Yes
Fit statistics
Observations 224 224 224 224
R2 0.906 0.903 0.904 0.906
Within R2 0.073 0.044 0.051 0.074
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
Table 3: Wage inequality regressions, wages ratio ($5^{th}/1^{st}$ dec.)
Dependent Variable: Ratio $5^{th}/1^{st}$ dec.
FE OLS
Model: (1) (2) (3) (4)
Variables
Between -4,368.036∗∗∗ -4,358.609∗∗∗
(1,552.729) (1,555.969)
Within -147.110 -62.273
(275.997) (290.821)
$\Delta C$607.034
GDPpc (log) -1.009 -1.003 -1.113 -1.023
(1.122) (1.173) (1.176) (1.126)
Population (log) -4.232 -4.258 -4.308 -4.239
(3.129) (3.226) (3.211) (3.144)
[…]
(0.011) (0.011) (0.011) (0.011)
[…]
(0.413) (0.420) (0.420) (0.418)
Fixed-effects
Country Yes Yes Yes Yes
Year Yes Yes Yes Yes
Fit statistics
Observations 224 224 224 224
R2 0.907 0.903 0.903 0.907
Within R2 0.090 0.048 0.055 0.090
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
Table 4: Wage inequality regressions, wages ratio ($9^{th}/5^{st}$ dec.)
Dependent Variable: Ratio $9^{th}/5^{th}$ dec.
FE OLS
Model: (1) (2) (3) (4)
Variables
Between -44.173 -39.238
(149.532) (147.483)
Within -33.360 -32.596
(70.242) (71.158)
$\Delta C$33.438
Population (log) 0.556 0.553 0.553 0.553
(0.502) (0.498) (0.498) (0.499)
GDPpc (log) -0.346∗∗ -0.353∗∗ -0.353∗∗ -0.353∗∗
(0.168) (0.162) (0.163) (0.162)
[…]
(0.002) (0.002) (0.002) (0.002)
[…]
(0.050) (0.048) (0.049) (0.049)
Fixed-effects
Country Yes Yes Yes Yes
Year Yes Yes Yes Yes
Fit statistics
Observations 224 224 224 224
R2 0.939 0.939 0.939 0.939
Within R2 0.109 0.110 0.110 0.110
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
Table 5: Labour share regressions
Dependent Variable: Labour share at constant 2015 USD
FE OLS
Model: (1) (2) (3) (4)
Variables
Between 10,825.259∗∗ 11,050.893∗∗
(4,288.268) (4,255.691)
Within -1,275.304 -1,490.403
(1,437.922) (1,428.899)
$\Delta C$ 99.792
(1,324.813)
Population (log) -34.886 -35.004 -34.852 -35.052
(18.884) (19.105) (19.209) (18.759)
GDPpc (log) -2.813 -3.202 -2.888 -3.150
(8.542) (8.738) (8.797) (8.688)
[…]
(0.095) (0.095) (0.096) (0.095)
[…]
(1.393) (1.407) (1.409) (1.424)
Fixed-effects
Country Yes Yes Yes Yes
Year Yes Yes Yes Yes
Fit statistics
Observations 224 224 224 224
R2 0.942 0.941 0.941 0.943
Within R2 0.300 0.287 0.285 0.302
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
5.3 Robustness checks
Given the small number of countries in the sample, we take preliminary steps to confirm that the results illustrated in the previous subsection are not driven by individual countries exhibiting extreme structural change and/or outcome values. In order to achieve this, we perform a set of leave-one-out regressions; for each outcome variable (employment growth, wage inequality and labour share) we run $N=30$ regressions, excluding in each one a different country. The results of this exercise are provided by figure 5, which shows the coefficient on the between component for each of these regressions. As shown in the figure, the coefficient remains always significant, indicating that excluding any country from the sample does not affect the results of any of the models estimated following equation 8.
Figure 5: Leave-one-out regressions results
Additionally, we are concerned that the signal captured by the structural change component (which is a composite measure combining changes in labour share within industries and their complexity) in explaining the outcomes of interest is driven by either of the elements that feed into the index. In order to test whether it is the case or not, we create a different structural change index, which uses the Shannon Entropy ($S$) instead of Complexity:
$LWS_{t}=\sum_{i=1}^{n}\theta_{i,t}S_{i,t};\qquad S=H(X)=-\sum_{i=1}^{n}p(x_{i})\,log\,p(x_{i})$ (12)
Table 6: Labour-weighted entropy regressions
[…]
FE OLS
Model: (1) (2) (3)
Variables
Between (Entropy) 5.773 60.682 87.949
(49.927) (37.579) (56.102)
Within (Entropy) -47.473 39.088 94.782
(88.854) (87.878) (127.908)
Population (log) -12.745 -1.189 -29.778
(11.211) (8.642) (20.182)
log(gdp_constant) -0.746 -3.862 -4.433
(4.060) (2.762) (8.187)
[…]
(0.058) (0.026) (0.097)
[…]
(0.952) (0.972) (1.330)
Fixed-effects
Country Yes Yes Yes
Year Yes Yes Yes
Fit statistics
Observations 224 224 224
R2 0.548 0.905 0.943
Within R2 0.134 0.064 0.304
Clustered (Country-Year) standard-errors in parentheses. Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
We proceed by regressing the newly created index – which is the between component of a different measure although related to LWF, named here Labour-Weighted Entropy – on employment growth, wage inequality, and labour share. Table 6 displays the results of this second robustness exercise, and indicates that the new measure of structural change does a very poor job in explaining the outcomes of interest. This results provides a further confirmation that the complexity-based structural change measure proposed in this paper is able to capture a strong signal that explains outcomes in employment creation and inequality, and proves to be a solid way to conceptualise structural change as the reallocation of workers across industries with different degrees of knowledge intensity.
6 Conclusions
The empirical analysis conducted in this paper has offered insights on the relationship between structural change – framed as the reallocation of labour towards more complex and knowledge intensive activities – and employment growth, wage inequality, and functional distribution of income. This relationship has been examined introducing several innovative methodological and analytical elements. In the first place, we have introduced a novel method to infer comparative advantages in an unbiased way, in what we believe to be the first empirical application of the BiWCM method proposed by Bruno et al. (2023). Secondly, we have constructed measures of Economic Fitness and Complexity that allow the comparison of these two indices over time (Mazzilli et al. 2024), overcoming a known limitation in the use of EC indices in econometric models (Caldarola et al. 2023). Third, we have introduced a new measure – the Labour-Weighted Fitness – whose variation allows to be decomposed in such a way as to capture the reallocation of labour across industries with different levels of complexity. This strategy has proved to have significant explanatory power with respect to the creation of employment, the distribution of wages, and the share of labour participation to the aggregate value added.
The preliminary results indicate that structural change towards complex industries slows down employment creation, due to the lower labour intensity of complex industries. However, the gap between top and bottom wages has shrunk, offering new insights on the inequality reduction brought about by structural change. It must be noted that wage inequality reduces as a result of an increase of the average wages at the first decile of the wage distribution, which in turn is the result of the loss of low-complexity (and likely low pay) jobs by the fittest most diversified countries. The increase in average wage also helps explain the relationship between structural change and the labour share, with the latter increasing when workers move to high-complexity industries. Once more, this effect is not likely to by ascribable to a higher rate of participation to the active workforce, but by the higher salaries paid to workers in countries specialised in complex industries. As a further step to confirm our results, we plan to examine the relationship between the structural change component of the LWF fitness and average wages, which may come in help to rationalise our findings on wage and functional distribution of income. Moreover, in order to confirm the robustness of the results, we are planning to devise an identification strategy based on a shift-share instrument, which appears at the moment the most suitable identification strategy given the nature of our (potentially) endogenous explanatory variable of interest.
Our results bear important consequences for policy making, as they highlight a trade-off between industrial upgrading and labour creation, while also dissecting the nature of the reduction of wage and functional inequality resulting from the movement of labour towards complex industries. Based on these results, policy-makers will have to device policy tools that steer structural change towards those industries that, at the same time, conflate job creation with higher technological sophistication, in order to sustain a process of inclusive economic growth. Such industries can be identified using the EFC framework with a focus on labour, rather than exports alone. Coupling the geography of productive capabilities with labour and inequality, enables us to advance with respect to the extant literature not only in providing a policy-relevant and data-driven toolkit to embrace socio-economic complexity, but also in offering a novel analytical and empirical understanding of a pathway not only geared towards boosting growth, but also to steering inclusive and sustainable development.
Appendix A Data reconstruction
The Eurostat’s Structural Business Survey data used to compute the employment based EFC measures used in this paper are one of the most comprehensive and detailed data sources of employment in Eurpean industries, classified at a fine level of disaggregation (NACE 4 digit). However, the dataset presents a large number of missing values, which hamper the construction of a specialisation matrix as explained in Section 4.1. In fact, the BiWCM and EFC algorithm require both a matrix with complete information.
In order to address this shortcoming in the SBS data, we have pursued, compared and validated seven different interpolation strategies, in order to retrieve a good approximation of the missing data. The first step has been to attempt the reconstruction of the SBS data following 7 alternative strategies, applied to the country-industry time series having a maximum 7 missing data points [^5] In the case of fully missing series, the country-industry couple has been excluded. In case of 8 out of 9 missing data points, the value has been assumed to be constant across the time series, extrapolating at constant values from the only available data point in the series.:
  1. 1.
    Internal missing values (those after the first available data point, and before the last) are interpolated using linear interpolation, while external missing values (before and after, respectively, the first and last data point available) have been extrapolated backwards taking the first available value as constant;
  2. 2.
    Internal missing values are interpolated using linear interpolation, while external missing values are extrapolated using the closest growth rate available, obtained by after interpolating internal missing values;
  3. 3.
    Internal missing values are interpolated using linear interpolation, while external missing values are extrapolated using average growth rate applied to the fist and last data points available after interpolating. This strategy produced negative values, which were constrained to 0;
  4. 4.
    Internal missing values are interpolated using linear interpolation, while external missing values are extrapolated using the first available growth rate by filling up/downwards available growth rates;
  5. 5.
    Internal missing values are interpolated using linear interpolation, while external missing values are extrapolated using moving average of growth rates, constructed using a rolling window of 3 years. The first two years are always ‘NA‘; second year is filled with moving average for t=2, and first year is just the growth with respect to the previous year;
  6. 6.
    Interpolate and extrapolate with linear fit computed using the available data points in each country-industry series;
  7. 7.
    Internal missing values are interpolated using linear interpolation, while external missing values are extrapolated using linear fit.
Each strategy is compared and validated using information from the complete country-industry time series. After filtering only the complete series from the dataset, we have created five different validation subsamples by randomising missing values in the complete dataset, imposing the following conditions:
  • Reflect NA frequency in original data (14% of data missing)
  • Reflect frequency in original data ± 5%,
  • Reflect frequency in original data + 10%
  • Impose 50% of missing values.
In order to pick the data reconstruction strategy that best approximates the observed data, we compare the Mean Absolute Error of each strategy across the 5 validation samples. As shown by table 7, the strategy that minimises the prediction error across the different validation subsamples is the first – i.e. the one that interpolates internal values linearly, and extrapolates external values taking the first and last available data point as constant, both backward and forward.
Table 7: Mean Absolute Errors
Reconstruction strategy
Subsample 1 2 3 4 5 6 7
NA 9% 174.96 189.05 179.47 192.23 189.44 222.55 197.73
NA 14% 214.95 227.60 215.10 230.20 227.45 273.22 240.86
NA 19% 254.61 294.14 267.06 297.80 292.47 323.64 286.27
NA 24% 311.79 329.90 315.13 331.72 325.46 395.41 355.26
NA 50% 692.86 788.71 749.16 796.20 784.38 819.18 777.39